# Electronic structure and transport in materials with flat bands: 2D materials and quasicrystals

### v1

@@ A.2.3 With defects
$\tau$ decreases when the temperature increases and/or static defect concentration increases. In actual 2D materials and aluminum-based quasicrystals at room temperature, realistic values of $\tau_{i}$ are a few $10^{-13}$ s [43, 246]. The conductivity $\sigma$ and diffusivity $\mathcal{D}$ can then be estimated from equations (63) and (64), where $\Delta X^{2}$ is the average square spreading in the crystal without defects. $\mathcal{D}$ is the sum of a Boltzmann contribution $\mathcal{D}_{\rm B}$ and a non-Boltzmann contribution $\mathcal{D}_{\rm NB}$:
@@ A.3 Real-space numerical method
This method makes it possible to calculate the quantum diffusion of charge carriers in supercell containing up to several $10^{7}$ orbitals. We can thus study the long-distance effects of non-periodic structures such as quasiperiodic structures, and/or introduce randomly distributed local structural static defects into the structure itself. The real-space method, which calculate $\Delta X^{2}(E,t)$ by using the polynomial expansion of the evolution operator on Tchebychev polynomials, has been developed by D. Mayou, S. N. Khanna, S. Roche, and F. Triozon [33, 34, 35, 36, 37]. A detailed presentation of the algorithm was given by F. Triozon in 2002 [428]. For a recent review see Ref. [366, 365]. It allows very efficient numerical calculations by recursion that take into account all quantum effects. It has been used for instance to study quantum transport in periodic approximants of quasicrystals [37, 62], disordered mono and multilayer graphene [82, 344, 85, 356, 345, 357, 346, 38, 358, 354, 125, 129, 133] and chemically doped graphene [83, 84], functionalized carbon nanotubes [362, 363, 364, 365], organic semiconductors [367, 368] and Perovskites [369], phosphorene [160], and other systems [366, 365].
Static defects are included directly in the structural modeling of the system and are randomly distributed on a supercell, which can contain several millions to tens of millions of atoms. This corresponds to typical sizes of about one micrometer square which allows to study systems with inelastic mean-free length of the order of few hundreds nanometers. Periodic boundary conditions are applied to this supercell. Note that for periodic systems with a large unit cell, such as quasicrystal approximants and twisted bilayer graphene, this supercell generally contains a large number of unit cells of the system under study.
Following the algorithm proposed in F. Triozon’s PhD thesis (2002) [428], the average square spreading $\Delta X^{2}$ is compute for any crystals taking into account the $s$, $p$ and $d$ orbitals in a Slater-Koster tight-binding (TB) model,
@@
where $\hat{U}(t)$ is the evolution operator at time $t$, $\delta$ is the Dirac function and ${\rm Tr}$ is the trace. Considering such a huge supercell, it is possible to evaluate the traces, ${\rm Tr}\,\hat{A}$, in the equation (78) by the average ${\langle\hat{A}\rangle}$ on a random phase state [37, 428], therefore,
@@
where $\{|i\rangle\}$ is the tight-binding basis, $N$ the number of orbitals in a supercell, and the real numbers $\theta_{i}$ are random. So the equation (78) becomes the ratio of two local densities of states (DOSs) on states $|\phi^{\prime}(t)\rangle=[\hat{X},\hat{U}(t)]|\phi\rangle$ and $|\phi\rangle$, respectively. For a given step time $T$, the evolution operator is decomposed as [37, 428],
$\hat{U}(T)={\rm e}^{-\frac{i}{\hbar}\hat{H}T}=\sum_{n=0}^{+\infty}c_{n}(T)\,Q_{n}(\hat{H})\approx\sum_{n=0}^{N_{Ch}}c_{n}(T)\,Q_{n}(\hat{H}),$ (80)
@@
As explained above, the calculation of $\Delta X^{2}(E,t)$ (Eq. (78)) is done by calculating the ratio of two DOSs. Theses DOSs calculations may be done by the recursion method (Lanczos algorithm) where the Hamiltonian is written as a tridiagonal matrix in real space [370] of dimension $N_{r}$. Usually we use $N_{r}$ between 1000 and 3000, and we checked that presented results do not change significantly when $N_{r}$ increases. The Lanczos method, leads to a convolution of the DOS by a Lorentzian function which a small width $\epsilon$. The DOS is thus obtained by a Lorentzian broadening of the spectrum and $\epsilon$ is a kind of energy resolution of the calculation. Usually we choose $\epsilon$ between 1 to 10 meV. For systems with a gap, to avoid the tail expansion of the Lorentzian function in the gap, it is more efficient to diagonalize the tridiagonal Hamiltonian matrix of dimension $N_{r}\times N_{r}$ and to compute the DOSs by a Gaussian broadening of the spectrum [369, 373]. Usually we have use a Gaussian standard deviation of $\sigma_{G}=5$ meV.

### v2

@@ A.2.3 With defects
$\tau$ decreases when the temperature increases and/or static defect concentration increases. In actual 2D materials and aluminum-based quasicrystals at room temperature, realistic values of $\tau_{i}$ are a few $10^{-13}$ s Mayou93 , Wu07 . The conductivity $\sigma$ and diffusivity $\mathcal{D}$ can then be estimated from equations (66) and (67), where $\Delta X^{2}$ is the average square spreading in the crystal without defects. $\mathcal{D}$ is the sum of a Boltzmann contribution $\mathcal{D}_{\rm B}$ and a non-Boltzmann contribution $\mathcal{D}_{\rm NB}$:
@@ A.3 Real-space numerical method
This method makes it possible to calculate the quantum diffusion of charge carriers in supercells containing up to several $10^{7}$ orbitals. We can thus study the long-distance effects of non-periodic structures such as quasiperiodic structures, and/or introduce randomly distributed local structural static defects into the structure itself. The real-space method, which calculates $\Delta X^{2}(E,t)$ by using the polynomial expansion of the evolution operator on Chebyshev polynomials, has been developed by D. Mayou, S. N. Khanna, S. Roche, and F. Triozon Mayou88 , Mayou95 , Roche97 , Roche99 , Triozon02 . A detailed presentation of the algorithm was given by F. Triozon in 2002 Triozon02_PhD . For a recent review see Refs. Fan14 , Fan20 . It allows very efficient numerical calculations by recursion that take into account all quantum effects. It has been used for instance to study quantum transport in periodic approximants of quasicrystals Triozon02 , Trambly17 , disordered mono and multilayer graphene Lherbier08 , Lherbier11 , Leconte11 , Leconte11b , Lherbier12 , Roche12 , Roche13 , Trambly13 , Missaoui17 , Missaoui18 , Omid20 , guerrero24_tBLG_disorder_Kubo , Guerrero25 and chemically doped graphene Lherbier08b , Leconte10 , functionalized carbon nanotubes Latil2004 , IshiiInelastic2010 , Jemai19 , Fan20 , organic semiconductors Fratini17 , Missaoui19 and perovskites Lacroix20 , phosphorene Jouda21 , and other systems Fan14 , Fan20 .
Static defects are included directly in the structural modeling of the system and are randomly distributed on a supercell, which can contain several millions to tens of millions of atoms. This corresponds to typical sizes of about one micrometer square which allows to study systems with inelastic mean free length of the order of few hundreds nanometers. Periodic boundary conditions are applied to this supercell. Note that for periodic systems with a large unit cell, such as quasicrystal approximants and twisted bilayer graphene, this supercell generally contains a large number of unit cells of the system under study.
Following the algorithm proposed in F. Triozon’s PhD thesis (2002) Triozon02_PhD , the average square spreading $\Delta X^{2}$ is computed for any crystal taking into account the $s$, $p$ and $d$ orbitals in a Slater-Koster tight-binding (TB) model,
@@
where $\hat{U}(t)$ is the evolution operator at time $t$, $\delta$ is the Dirac function and $\rm Tr$} is the trace. Considering such a huge supercell, it is possible to evaluate the traces, $^{\rm Tr}\,\hat{A}$, in the equation (82) by the average $^{\langle\hat{A}\rangle}$ on a random phase state Triozon02 , Triozon02_PhD , therefore,
@@
where $\{|i\rangle\}$ is the tight-binding basis, $N$ the number of orbitals in a supercell, and the real numbers $\theta_{i}$ are random. So the equation (82) becomes the ratio of two LDOSs on states $|\phi^{\prime}(t)\rangle=[\hat{X},\hat{U}(t)]|\phi\rangle$ and $|\phi\rangle$, respectively. For a given step time $T$, the evolution operator is decomposed as Triozon02 , Triozon02_PhD ,
@@
[…]
As explained above, the calculation of $\Delta X^{2}(E,t)$ (Eq. (82)) is done by calculating the ratio of two DOSs. These DOS calculations may be performed by the recursion method where the Hamiltonian is first written as a tridiagonal matrix in real space of dimension $N_{r}$ using the Lanczos algorithm Lanczos50 , Cullum1984 , Pettifor , Dagotto1994 . Usually we use $N_{r}$ between 1000 and 3000, and we checked that results do not change significantly when $N_{r}$ increases. One can then use the Lanczos coefficients for a continued-fraction expansion of the DOS Haydock1972 , Pettifor , Gagliano1987 , Dagotto1994 , Schenk08 . This recursion method leads to a convolution of the DOS by a Lorentzian function which a small width $\epsilon$. The DOS is thus obtained by a Lorentzian broadening of the spectrum and $\epsilon$ is a kind of energy resolution of the calculation. Usually we choose $\epsilon$ between 1 and 10 meV. Alternatively, one can also diagonalize the tridiagonal Lanczos matrix of dimension $N_{r}\times N_{r}$ and apply a suitable broadening once the poles and their weights are known from the diagonalization. In the case of Lorentzian broadening, both approaches are equivalent Gagliano1987 , Dagotto1994 . However, to avoid the tail expansion of the Lorentzian function in the gap for systems with a gap, it is preferable to diagonalize the tridiagonal Lanczos matrix and to evaluate the DOS by a Gaussian broadening of the spectrum Lacroix20 , Lacroix_these . Usually we have used a Gaussian standard deviation of $\sigma_{G}=5$ meV.